### v1

# Bayesian model selection and misspecification testing in imaging inverse problems only from noisy and partial measurements

Abstract
Modern imaging techniques heavily rely on Bayesian statistical models to address difficult image reconstruction and restoration tasks. This paper addresses the objective evaluation of such models in settings where ground truth is unavailable, with a focus on model selection and misspecification diagnosis. Existing unsupervised model evaluation methods are often unsuitable for computational imaging due to their high computational cost and incompatibility with modern image priors defined implicitly via machine learning models. We herein propose a general methodology for unsupervised model selection and misspecification detection in Bayesian imaging sciences, based on a novel combination of Bayesian cross-validation and data fission, a randomized measurement splitting technique. The approach is compatible with any Bayesian imaging sampler, including diffusion and plug-and-play samplers. We demonstrate the methodology through experiments involving various scoring rules and types of model misspecification, where we achieve excellent selection and detection accuracy with a low computational cost. [^1] The code used to run the experiments is publicly available at https://github.com/aleph-group/Priors_selection.
1 Introduction
Preliminaries
Modern quantitative and scientific imaging techniques heavily rely on statistical models and inference methods to analyze raw sensor data, reconstruct high-quality images, and extract meaningful information [4]. Despite the diversity of imaging modalities and applications, most statistical imaging methods aim to infer an unknown image $x_{\star}\in\mathbb{R}^{n}$, from a measurement $y\in\mathbb{R}^{m}$, modeled as a realization of
$\mathbf{y}\sim P(A(x_{\star}))$ (1)
where $A$ is an experiment-specific measurement operator representing deterministic physical aspects of the sensing process, and $P$ is a statistical noise model [21]. Common examples include image denoising, demosaicing, deblurring, and tomographic reconstruction [4].
A key common feature across statistical imaging is that recovering $x_{\star}$ from $y$ involves solving an inverse problem that is not well-posed, requiring regularization to stabilize the inversion. The Bayesian imaging paradigm addresses regularization by treating $x_{\star}$ as a random variable $\mathbf{x}$ and incorporating prior knowledge through the marginal $p(x)$. This prior is then combined with the likelihood function $p(y|x)$ by using Bayes’ theorem to obtain the posterior distribution
$p(x|y)=\frac{p(y|x)p(x)}{\int p(y|\tilde{x})p(\tilde{x})\textrm{d}\tilde{x}}\,$,
which underpins all inferences about $\mathbf{x}$ having observed $\mathbf{y}=y$ [43]. Beyond producing estimators, modern Bayesian imaging methods increasingly quantify uncertainty in the reconstruction, an essential component for reliable interpretation and robust integration with decision-making processes. Of course, modeling choices may strongly influence the delivered inferences, making the development of ever more accurate Bayesian imaging models a continual focus of research.
Modern Bayesian imaging methods increasingly use highly informative image priors encoded by deep learning models that deliver unprecedented estimation accuracy [18, 38]. Notable examples of Bayesian imaging frameworks with data-driven priors include plug-and-play Langevin samplers [27, 42, 25], denoising diffusion models [55, 11, 46, 37], distilled diffusion models [47, 34], flow matching [32], and conditional GANs [2, 3]. In addition, while traditional approaches to developing data-driven image priors required large amounts of clean training data, modern methods increasingly learn image models directly from measurement data [7]. These models can also be designed to exhibit the mathematical regularity needed for integration into optimization algorithms and Bayesian sampling machinery [39].
However, while already widely deployed in photographic imaging pipelines, leveraging data-driven priors for quantitative and scientific imaging remains challenging due to the stricter requirements for reliability and accuracy. For example, data-driven priors can lead to strongly biased inferences if, during deployment, the encountered image $x_{\star}$ is poorly represented in the training data. In such cases, highly informative priors may override the likelihood $p(y|x)$, particularly in ill-posed or ill-conditioned problems where the likelihood has poor identifiability. It is therefore essential to equip critical imaging pipelines with the ability to self-diagnose model misspecification. Similarly, multiple data-driven priors and likelihoods may be available for inference, each reflecting different assumptions about the sensing process and the scene; assumptions that are often unverifiable in practice. Hence, robust imaging pipelines must be able to objectively compare alternative models based solely on measurement data.
Problem statement
This paper considers the problem of objectively comparing and diagnosing misspecification in Bayesian imaging models directly from measurement data, without access to ground truth. The focus is on modern data-driven image priors encoded by large machine learning models, which are highly informative and may be improper.
Contributions
We herein propose a statistical methodology for performing Bayesian model selection and misspecification diagnosis in large-scale imaging inverse problems. Our proposed methodology is fully unsupervised, in that the analyses solely use a single noisy measurement $y$. This is achieved by leveraging measurement splitting by noise injection [40, 36], also known as data fission [29], in order to construct a self-supervising Bayesian cross-validation procedure. The methodology is agnostic to the class of image priors used and fully compatible with modern priors encoded by deep learning models. In addition, the method is computationally efficient and can be straightforwardly integrated within widely used Bayesian imaging sampling strategies, such as Langevin and guided denoising diffusion samplers. We demonstrate the effectiveness of our approach through numerical experiments related to image deblurring with plug-and-play Langevin samplers and denoising diffusion models for photographic and magnetic resonance images, where we report excellent model selection and misspecification detection accuracy even in challenging settings.
2 Background
Prior predictive checking
evaluates the model $p(x|y)$ by comparing the observation $y$ to predictions of $\mathbf{y}$ derived from the model [14]. Such checks often use the prior predictive distribution, with density $p(y)=\int p(y|x)p(x)\textrm{d}x$, or more generally an expected utility loss $\Phi(y)=\int\phi(y,x)p(x)\textrm{d}x$, where $\phi(y,x)$ quantifies the discrepancy between a possible $x$ and $y$. Prior predictive checks implicitly view $p(\mathbf{x},\mathbf{y})$ as a generative model for $(\mathbf{x},\mathbf{y})$ and they provide a useful lens to examine the implications of specific prior and likelihood choices. However, they do not evaluate how well the model supports inference on $\mathbf{x}$ after fitting to $\mathbf{y}=y$, nor do they reveal how specific forms of model misspecification affect particular inferences. Additionally, prior predictive checks are not well-defined when $p(\mathbf{x})$ is improper even if the resulting posterior $p(\mathbf{x}|\mathbf{y}=y)$ is well-posed and yields meaningful accurate inferences, as is often the case in Bayesian imaging models.
Posterior predictive checking
evaluates the model $p(\mathbf{x},\mathbf{y})$ through the prediction of a new measurement $\mathbf{y}^{+}\sim P(Ax_{\star})$ stemming from a hypothetical experiment replication, conditionally to $\mathbf{y}=y$ [14]. Such checks leverage the posterior predictive distribution, with density $p(y^{+}\mid y)=\int p(y^{+}|x)p(x|y)\textrm{d}x$ where the unknown image $\mathbf{x}$ is drawn from $p(x|y)$. Posterior predictive checks reveal model misfit by identifying discrepancies between the prediction $\mathbf{y}^{+}|\mathbf{y}=y$ derived from $p(\mathbf{x}|\mathbf{y}=y)$ and the observed measurement $\mathbf{y}=y$. Again, both application-agnostic and task-specific scoring rules can be used to probe targeted aspects of the model. However, posterior checks are often overly optimistic, as predictions are conditioned on the observed data and thus biased towards agreeing with it [14].
Bayesian cross validation
is a powerful partial posterior predictive approach that mitigates the bias of conventional posterior predictive checks by holding out part of the data, fitting the model to the remainder, and evaluating predictive performance on the held-out set. This yields more reliable diagnostics, as it breaks the circularity of using the same data for both model fitting and evaluation [49, 15, 10]. To make full use of the data, cross-validation employs randomization, repeatedly fitting and evaluating across multiple data partitions. While widely adopted in other domains, Bayesian cross-validation remains largely unexplored in computational imaging, where typically only a single measurement is available. Unfortunately, obtaining two independent measurements of the same scene is often not possible, as imaging experiments occur under conditions that are ephemeral due to dynamic scenes, non-static sensors, and operational constraints.
Unsupervised Bayesian model selection
uses strategies similar to model evaluation -namely prior, posterior or partial predictive summaries- but differs fundamentally in its purpose. Model selection aims to rank competing models and identify the one that best explains the observed data, rather than assessing individual model adequacy. Unsupervised Bayesian model selection for computational imaging often relies on the (prior predictive) marginal likelihood $p(y)=\mathbb{E}(p(y\mid\mathbf{x}))=\int p(y,{x})\textrm{d}x$, particularly through the use of so-called Bayes factors to assess the relative fit-to-data of competing models. However, computing marginal likelihoods for image data is notoriously challenging due to the high dimensionality involved. Early approaches have used harmonic mean estimators [12], while recent efforts have employed nested samplers specifically designed for this task [45, 6, 35]; however, these remain computationally expensive and difficult to scale. Approximations based on empirical Bayesian residuals [51] offer a tractable alternative, but their reliability is limited [33]. One can also consider supervised Bayesian model selection, relying on reference images and controlled experiments. However, this approach is impractical in many application domains where acquiring reliably representative reference data is infeasible.
Out-of-distribution detection.
In Bayesian imaging, out-of-distribution detection (OOD) methods are predominantly used to identify situations of prior misspecification with respect to datasets. As stated previously, this is especially important when using highly informative priors encoded by large machine learning models. Several supervised OOD methods have been recently proposed in the literature [30, 17, 54, 13], along with a recent unsupervised OOD method specifically designed for diffusion models [44]. To the best of our knowledge, no existing methods can diagnose OOD based on a single measurement or address general Bayesian imaging reconstruction techniques.
3 Proposed method
3.1 Bayesian cross-validation by data fission
We now present our methodology for Bayesian model selection and misspecification testing. Suppose for now the availability of two independent measurements $\mathbf{y}^{+},\mathbf{y}^{-}\sim P(A(x_{\star}))$ from replication of the experiment. Adopting a partial predictive approach, we evaluate a model $\mathcal{M}$, comprising a prior and likelihood, by computing a summary of the form [49]
$\Psi$ $(\mathcal{M})=\mathbb{E}_{\mathbf{y}^{+},\mathbf{y}^{-}}\left[S(p_{\mathcal{M}}(\mathbf{y}^{+}|\mathbf{y}^{-},y^{+})\right]\,$, (2)
$=\int S(p_{\mathcal{M}}(\mathbf{y}^{+}|\mathbf{y}=y^{-}),y^{+})p_{\mathcal{M}}(y^{-},y^{+}){\rm d}y^{-}\,{\rm d}y^{+}\,$,
where $S:\mathcal{P}\times\mathbb{R}^{m}\mapsto\mathbb{R}_{+}$ is a scoring rule [16] that takes a predictive density $p\in\mathcal{P}$, with $\mathcal{P}$ being a probability measure, and a realization mapping it to a numerical assessment of that prediction. In the case of (2), we summarize the models’ capacity to predict $\mathbf{y}^{+}$ having observed $\mathbf{y}^{-}$, under the assumptions encoded by $p(y^{-},y^{+})=\int p(y^{+}|x)p(y^{-}|x)p(x)\textrm{d}x$ as described by model $\mathcal{M}$. With regards to $S$, a classic choice is the logarithmic rule ${S}(p(\mathbf{y}^{+}|y^{-}),{y}^{+})\coloneqq\log p(y^{+}|y^{-})=\log\int p(y^{+}|x)p(x|y^{-})\textrm{d}x$, which is known to be strictly proper [16]. Other rules allow probing of $\mathcal{M}$ for particular forms of misspecification; examples tailored for imaging are provided later.
Summaries of the form (2) are usually computed approximately by cross-validation, with K-fold randomization of the data partition. However, implementing Bayesian cross-validation in imaging is challenging, as often only a single data point $y$ is available. To overcome this fundamental difficulty, our approach leverages data fission [29], a form of measurement splitting by noise injection used in computer vision [40, 36]. This leads to a Bayesian cross-validation approach that, from a single measurement $y$, offers a trade-off between accuracy and computational efficiency.
Measurement splitting strategies partition a single observed outcome $\mathbf{y}=y$ from $\mathbf{y}\sim P(A(x_{\star}))$ into two synthetic measurements $\mathbf{y}^{+}$ and $\mathbf{y}^{-}$ that are conditionally independent given $x_{\star}$. For presentation clarity, we introduce this step for problems involving additive Gaussian noise, and subsequently extend the approach to other noise models. Suppose that $\mathbf{y}\sim\mathcal{N}(A(x_{\star}),\Sigma)$ and let $\mathbf{w}\sim\mathcal{N}(0,\Sigma)$. Then, for any $\alpha\in(0,1)$,
[…]
with $c_{\alpha}=\sqrt{\alpha/(1-\alpha)}$ are independent Gaussian variables conditionally to $x_{\star}$, with mean $A(x_{\star})$ and covariance proportional to $\Sigma$. For the specific case of $\alpha=0.5$, they are i.i.d. with marginal distribution $\mathcal{N}(A(x_{\star}),2\Sigma)$. For $\alpha\neq 0.5$, we have that the information in $\mathbf{y}$ is divided unequally between $\mathbf{y}^{+}$ and $\mathbf{y}^{-}$; reducing $\alpha$ brings $\mathbf{y}^{+}$ closer to $\mathbf{y}$ and reduces the correlation between $\mathbf{y}$ and $\mathbf{y}^{-}$. Equivalent splitting strategies are available for other noise models from the natural exponential family [36], including Poisson and Gamma noise commonly encountered in imaging.
By combining (2) with measurement splitting, our proposed Bayesian cross-validation approach evaluates a model $p_{\mathcal{M}}(x,y)$ through its capacity to deliver accurate predictions of $f^{+}_{\alpha}(y,\mathbf{w})$ from $f^{-}_{\alpha}(y,\mathbf{w})$; i.e.,
$\Phi(\mathcal{M})=\mathbb{E}_{\mathbf{w}}\left[\mathbb{E}_{\mathbf{x}\mid f^{-}_{\alpha}(y,\mathbf{w}),\mathcal{M}}\left[\phi_{\mathcal{M}}(f^{+}_{\alpha}(y,\mathbf{w}),\mathbf{x}))\right]\right]$ (4)
$=\int\phi_{\mathcal{M}}(f^{+}_{\alpha}(y,{w}),{x})\,p_{\mathcal{M}}({x}\mid f^{-}_{\alpha}(y,{w}))p({w})\textrm{d}{x}\textrm{d}{w}$
where $\phi_{\mathcal{M}}:\mathbb{R}^{m}\times\mathbb{R}^{n}\mapsto\mathbb{R}_{+}$ quantifies the discrepancy between a possible $x$ and $y^{+}$, leading to a scoring rule $\mathbb{E}_{\mathbf{x}\mid f^{-}_{\alpha}(y,\mathbf{w}),\mathcal{M}}\left[\phi_{\mathcal{M}}(f^{+}_{\alpha}(y,\mathbf{w}),\mathbf{x}))\right]$ for the prediction of $\mathbf{y}^{+}$ from $y^{-}$ (related to each other via $\mathbf{x}\sim p_{\mathcal{M}}(x|y^{-})$, which is marginalized out). The expectation over the noise $\mathbf{w}$ plays a role analogous to randomized data partitions in K-fold cross-validation, with $\alpha$ controlling the share of information in $y$ that is held out.
3.2 Scoring rules for probing imaging models
Below, we discuss two specific scoring rules we recommend for imaging applications.
Likelihood-based rule
To probe the likelihood $p(y|x)$, we use a rule based on the log likelihood $\phi^{1}_{\mathcal{M}}(f^{+}_{\alpha}(y,\mathbf{w}),\mathbf{x})=\log p_{\mathcal{M}}(f^{+}_{\alpha}(y,\mathbf{w})|\mathbf{x})$ and obtain
$\Phi^{1}(\mathcal{M})=\mathbb{E}_{\mathbf{w}}\left[\mathbb{E}_{\mathbf{x}\mid f^{-}_{\alpha}(y,\mathbf{w}),\mathcal{M}}\left[\log p_{\mathcal{M}}(f^{+}_{\alpha}(y,\mathbf{w})|\mathbf{x})\right]\right]\,.$ (5)
This rule is closely related to the logarithmic score applied to (2) via Jensen’s inequality [20]. However, we recommend it over the logarithmic score due to its significantly greater numerical stability [5].
Posterior-based rule
Consider a severely ill-posed inverse problem where $A$ is severely rank deficient and therefore the observations are not very informative. In that case, the rule based on the log likelihood will have poor discrimination w.r.t. the properties of the prior. For example, in the case of a linear Gaussian model, $\log p_{\mathcal{M}}(f^{+}_{\alpha}(y,{w})|\mathbf{x})\propto\|f^{+}_{\alpha}(y,{w})-A\mathbf{x}\|_{2}^{2}$ will not be sensitive to information about $p_{\mathcal{M}}(x|f^{-}_{\alpha}(y,\mathbf{w}))$ in the null space of A. In this scenario, we recommend using a rule that incorporates $p_{\mathcal{M}}(x|f^{+}_{\alpha}(y,\mathbf{w}))$, so that there is a direct comparison between $p_{\mathcal{M}}(x|f^{+}_{\alpha}(y,\mathbf{w}))$ and $p_{\mathcal{M}}(x|f^{-}_{\alpha}(y,\mathbf{w}))$ without the action of $A$. For example,
$\phi^{2}_{\mathcal{M}}(f^{+}_{\alpha}(y,\mathbf{w}),\mathbf{x})=\mathbb{E}_{\mathbf{x}^{\prime}\mid f^{+}_{\alpha}(y,\mathbf{w}),\mathcal{M}}\left[s_{\rho}(\mathbf{x},\mathbf{x}^{\prime})\right],$ (6)
where $s_{\rho}:\mathbb{R}^{k}\times\mathbb{R}^{k}\mapsto\mathbb{R}_{+}$ is a discrepancy in an embedding space tailored for a particular task, and is generated by the map $\rho:\mathbb{R}^{n}\mapsto\mathbb{R}^{k}$. The resulting summary reads
$\Phi^{2}_{y}(\mathcal{M})=\mathbb{E}_{\mathbf{w}}\left[\mathbb{E}_{\mathbf{x}\mid f^{-}_{\alpha}(y,\mathbf{w}),\mathcal{M}}\left[\mathbb{E}_{\mathbf{x}^{\prime}\mid f^{+}_{\alpha}(y,\mathbf{w}),\mathcal{M}}\left[s_{\rho}(\mathbf{x},\mathbf{x}^{\prime})\right]\right]\right].$ (7)
A standard choice for the discrepancy would be $s_{\rho}({x},{x}^{\prime})=\|\rho(x)-\rho(x^{\prime})\|_{2}$. Depending on the characteristics of the inverse problem and the model, different embedding spaces can be considered. For a distortion-focused comparison, the embedding mapping $\rho(\cdot)$ would be the identity. However, we can use LPIPS-based embedding [54] for a perception-focused comparison, or CLIP-based embedding [41] for a semantic-focused comparison.
Monte Carlo approximation
In practice, we approximate the expectations in the comparison metrics using Monte Carlo sampling. For the likelihood-based metric under Gaussian noise with a diagonal covariance matrix, we compute the negative log-likelihood (omitting the normalization constant), as follows:
$\widehat{\Phi}^{1}_{y}(\mathcal{M})=\frac{1}{KN}\sum_{k=1}^{K}\sum_{n=1}^{N}\|y+c_{\alpha}w_{k}-A(x_{k,n})\|_{2}^{2},$ (8)
where $x_{k,n}$ follows the posterior $(\mathbf{x}^{-}\mid f^{-}_{\alpha}(y,{w}_{k}),\mathcal{M})$ and ${w}_{k}$ is a realization of $\mathcal{N}(0,\sigma I_{m})$. For the posterior-based rule with an LPIPS embedding $\rho_{\rm L}$, we have
$\widehat{\Phi}^{2}_{y}(\mathcal{M})=\frac{1}{KNL}\sum_{k,n,l=1}^{K,N,L}\|\rho_{\rm L}(x^{-}_{k,n})-\rho_{\rm L}(x^{+}_{k,l})\|_{2},$ (9)
where $x^{-}_{k,n}$ and $x^{+}_{k,l}$ are respectively samples from $(\mathbf{x}^{-}\mid f^{-}_{\alpha}(y,{w}_{k}),\mathcal{M})$ and $(\mathbf{x}^{+}\mid f^{+}_{\alpha}(y,{w}_{k}),\mathcal{M})$. Our experiments suggest that the estimators $\widehat{\Phi}^{1}_{y}(\mathcal{M})$ and $\widehat{\Phi}^{2}_{y}(\mathcal{M})$ are accurate even with few samples.
3.3 Relation with posterior predictive checks and the marginal likelihood
Let us now consider a single splitting noise realization $\mathbf{w}=w$. For ease of presentation we note $y^{+}=f^{+}_{\alpha}(y,{w})$ and $y^{-}=f^{-}_{\alpha}(y,{w})$ and use the splitting from (3). If we choose the likelihood $\phi^{3}_{\mathcal{M}}(y^{+},\mathbf{x})=p_{\mathcal{M}}(y^{+}|\mathbf{x})$, the proposed metric in Eq. (4) reads
$\begin{split}\Phi^{3}_{y}(\mathcal{M})&=\mathbb{E}_{\mathbf{x}\mid y^{-},\mathcal{M}}\left[p_{\mathcal{M}}(y^{+}|\mathbf{x})\right]=p_{\mathcal{M}}(y^{+}|y^{-})\\ &=\int p_{\mathcal{M}}(y^{+}|{x})p_{\mathcal{M}}(x|y^{-}){\rm d}x\end{split}$ (10)
which is the posterior predictive check for model $\mathcal{M}$ on the “new” observation $y^{+}$ conditioned to the previous observation $y^{-}$. In the limit of $\alpha$ tending to zero, we have that $y^{+}$ tends to $y$ and $y^{-}$ to an independent noise realization. Hence, for $\lim_{\alpha\to 0}\Phi^{3}_{y}(\mathcal{M})$ we obtain
$\mathbb{E}_{\mathbf{x}\mid\mathcal{M}}\left[p_{\mathcal{M}}(y|\mathbf{x})\right]=p_{\mathcal{M}}(y)=\int p_{\mathcal{M}}(y|{x})p_{\mathcal{M}}(x){\rm d}x,$ (11)
which is the marginal likelihood. The main difference between the two previous formulations is that, in the first one, $\mathbf{x}$ follows the partial posterior $p(x|y^{-},\mathcal{M})$ instead of the prior $p(x|\mathcal{M})$. Conditioning on the variable $y^{-}$, a noisier version of $y$, greatly helps to improve the behavior of the estimator. Each sample from the pseudo posterior is more likely to have a higher likelihood value and to contribute to the calculation of the expectation. We approximate the estimator (11) as
$\widehat{p}_{\mathcal{M}}(y^{+}|y^{-})=\frac{1}{KN}\sum_{k=1}^{K}\sum_{n=1}^{N}p_{\mathcal{M}}(y+c_{\alpha}w_{k}|x_{k,n}),$ (12)
where $x_{k,n}$ follows the posterior $\mathbf{x}|y-w_{k}/c_{\alpha},\mathcal{M}$ and $w_{k}$ is a realization of $\mathcal{N}(0,\sigma^{2}I_{m})$.
The role of $\alpha$ is to control the split of information between the conditioning variable $y^{-}$, helping to ease the evidence calculation, and the estimator variable $y^{+}$, which we use to compute the marginal likelihood and evaluate model fit-to-data.
4 Experimental results
4.1 Error analysis in the Gaussian case
We first study a toy Gaussian model, designed to illustrate the proposed methodology under various degrees of model misspecification, model size, and splitting parameter $\alpha$. We assume that $\mathbf{y}=\mathbf{x}+\mathbf{e}$, where $\mathbf{e}\sim\mathcal{N}(0,\sigma^{2}I_{m})$ and $\mathbf{x}\sim\mathcal{N}(0,\sigma^{2}_{x}I_{m})$ are independent of $\mathbf{e}$. For ease of presentation, we use the notation $y^{+}$ and $y^{-}$ from Section 3.3 . For this model, we have a Gaussian posterior $p(x|y^{-})=p_{\mathcal{N}}(x|\frac{\alpha}{\alpha\sigma_{x}^{2}+\sigma^{2}}y^{-},\frac{\sigma^{2}\sigma_{x}^{2}}{\sigma^{2}+\alpha\sigma_{x}^{2}}I_{m})$, where the predictive density $p(y^{+}|y^{-})$ is tractable (see Section 1 of the supplementary material (SM)).
We draw realizations from $\mathbf{y}$ with $\sigma^{2}=1$ and posit that $\mathbf{x}\sim\mathcal{N}(0,{\sigma^{\prime}}^{2}_{x}I_{m})$ to study the impact of misspecification. Fig. 1 shows the expectation of the marginal log-likelihood ratio $\log({p(\mathbf{y}^{+}|\mathbf{y}^{-},\sigma^{2}_{x})}/{p(\mathbf{y}^{+}|\mathbf{y}^{-},\sigma_{x}^{\prime 2})})$ as a function of $\sigma_{x}^{\prime}$ for different values of $\alpha$, as estimated by averaging over $K=250$ realizations of $\mathbf{w}$ and when $m=1000$. We observe that, as expected, model discrimination improves as $\alpha$ decreases and more information is held out in $\mathbf{y}^{+}$ for model evaluation (recall that $\alpha\rightarrow 0$ leads to the marginal likelihood, which is excellent for model discrimination but often computationally intractable). Moreover, we see in Fig. 2 that averaging $K$ realizations of $\mathbf{w}$ reduces the bias introduced by measurement splitting, similarly to randomization in K-fold cross-validation. With regards to computational cost, reducing $\alpha$ increases the number of Monte Carlo samples required to reliably approximate $p(y^{+}|y^{-})$, highlighting a trade-off between evaluation accuracy and efficiency (see SM, Section 1).
Figure 1: Log difference between $p(y^{+}|y^{-},\sigma_{x}^{2})$ and $p(y^{+}|y^{-},\sigma_{x}^{\prime 2})$ as a function of $\sigma_{x}^{\prime}$ and for different $\alpha$, averaged over the injected noise $\mathbf{w}$. The true prior standard deviation is $\sigma_{x}=1$.
Figure 2: Log difference between $p(y^{+}|y^{-},\sigma_{x}^{2})$ and $p(y^{+}|y^{-},\sigma_{x}^{\prime 2})$ as a function of $\sigma_{x}^{\prime}$ and for different numbers of noise realizations $K$, with $\alpha=0.5$. The true prior standard deviation is $\sigma_{x}=1$.
4.2 Unsupervised likelihood model selection
We now consider an image deblurring problem $\mathbf{y}=Ax_{\star}+\mathbf{e}$, where $\mathbf{e}\sim\mathcal{N}(0,\sigma^{2}I_{m})$ with $\sigma=0.1$ and where $A$ is a circulant blur operator implementing the action of a blurring kernel $\kappa_{\operatorname{GT}}$. Given blur kernels from the Moffat, Laplace, Uniform, and Gaussian parameter families presented in Fig. 4, set to be as close as possible, we wish to identify the ground truth kernel relating a measurement $y$ to $x_{\star}$. Refer to the SM, Section 2, for the parametric kernel forms.
For each test image in Fig. 3, of size $256\times 256$ pixels, we generate 5 noisy measurements using the 5 kernel as ground truth. We then compute the value of the log-likelihood-based estimator $\widehat{\Phi}^{1}$ for each observation and each one of the considered blur kernels, seeking to use $\widehat{\Phi}^{1}$ to identify the correct kernel. We adopt a Langevin PnP approach [28] and use the gradient step denoiser [19] as prior together with the SK-ROCK algorithm [1] for posterior sampling. To compute $\widehat{\Phi}^{1}_{y}$, we set $\alpha=0.5$ and draw $K=10$ realizations of $\mathbf{w}$ and $N=100$ posterior samples per realization. Tab. 1 reports the model selection accuracy for our method when each observation is analyzed separately (single shot), and when we assume that the blur kernel is shared across the three images (few shot). We observe that our method correctly identifies the blur kernel from a single measurement in over $85\%$ of the cases, and with perfect accuracy when pooling three measurements. For comparison, we also report the Bayesian residual method [51] and the empirical Bayesian variant that improves model selection performance by automatically calibrating model parameters [50]. Their accuracy is noticeably lower, in the order of $40\%$ to $60\%$. Please refer to SM, Section 2, for implementation details.
Figure 3: Examples of blurred measurements, generated by using the blur kernel $\kappa_{\mathcal{G}}(2)$.
Single Shot Few Shot
[…]
Bayes Res. [51] $\% 40.0\%$
[…]
Table 1: Accuracy of likelihood model selection, using the the proposed summary $\widehat{\Phi}^{1}$ and two variants of the baseline method [51], from a single measurement (single shot) or three measurements (few shot).
Figure 4: Profile of the considered blur kernels, their similarity makes model selection difficult.
4.3 Prior selection and OOD detection
We now explore our estimator’s ability to objectively compare different image priors and diagnose prior misspecification in OOD situations. We focus on priors represented by denoising diffusion models and use the DiffPIR algorithm [55] for posterior sampling.
4.3.1 Deblurring of natural images
Figure 5: Posterior samples from $p(x|y^{-})$ for $\alpha=0.1$, $\sigma_{\kappa}=0.5$, for some test natural image examples.
We first consider a deblurring problem on natural images of size $256\times 256$ pixels. We use two Diffusion UNet models from Choi et al. [8] as priors, which were trained on color images from FFHQ and AFHQ-dogs respectively. We define the forward operator $A$ as an isotropic Gaussian blurring operator with bandwidth $\sigma_{\kappa}\in\{0.05,\;2,\;5\}$ to reflect mild, moderate and high blur. Two datasets are defined: a reference in distribution (ID) subset of $60$ images from FFHQ [24], and test dataset composed of $90$ images from AFHQ [9], CelebA-HQ [22], LSUN-Bedrooms [52], Met-Faces [23], CBSD68 [31], and FFHQ, representing different degrees of prior misspecification. Indeed, while Bedrooms, CBSD68 and AFHQ images are strongly OOD, the images from Met are only moderately OOD and constitute a limit case. Celeb images stem from a different dataset but should be considered ID.
We compute the estimators $\widehat{\Phi}^{1}_{y}$ and $\widehat{\Phi}^{2}_{y}$ for the reference images and test images, using $K=10$ noise realizations with $N=20$ samples each, and $\alpha=0.1$. Fig. 7 depicts the values of the estimators $\widehat{\Phi}^{1}_{y}$ and $\widehat{\Phi}^{2}_{y}$ on the reference dataset and the test datasets. We observe that $\widehat{\Phi}^{2}_{y}$ is highly sensitive to OOD situations, whereas $\widehat{\Phi}^{1}_{y}$ has more limited value in this case.
For OOD detection, we consider the null hypothesis to be “in distribution”, and we define a simple statistical test by setting a threshold at the $95$-th percentile of $\widehat{\Phi}^{2}_{y}$ over the reference dataset. Tab. 2 reports the Type I error probability and power for each test subset, at significance level $5\%$. Observe that testing with $\widehat{\Phi}^{2}_{y}$ achieves a Type I error close to the desired $5\%$ on the two ID datasets, and excellent power on the moderate and strongly OOD datasets. As expected, the power of the test decreases as the blur strength $\sigma_{\kappa}$ increases and removes fine detail, especially in mild OOD cases.
The effectiveness of $\widehat{\Phi}^{2}_{y}$ stems from the fact that, when $x_{\star}$ is OOD and $\alpha$ is small, the noise imbalance between $y^{+}$ and $y^{-}$ creates a noticeable perceptual discrepancy between the posterior samples from $p(x|y^{+})$ and $p(x|y^{-})$. To illustrate this, Figure 6 depicts samples from the posterior distributions $p(x|y^{+})$ and $p(x|y^{-})$ under both well-specified and strongly misspecified priors. As the blur strength increases, perceptual hallucinations become more pronounced in the OOD model’s reconstructions. This effect persists, though more weakly, under mildly misspecified priors, resulting in a drop in detection power at high blur levels. To illustrate a mild OOD situation, Figure 8 shows a Met-Faces example that is correctly identified as OOD for $\sigma_{\kappa}=0.5$ and $\sigma_{\kappa}=2$, but misclassified for $\sigma_{\kappa}=5$.
$\sigma_{\kappa}=0.5$ $\sigma_{\kappa}=2$ $\sigma_{\kappa}=5$
Type I Error
FFHQ 0% 6.7% 6.7%
Celeb 6.7% 6.7% 6.7%
Power
Moderate OOD 86.7% 73.3% 60%
Strong OOD 100% 100% 100%
Table 2: Type I error rate (incorrect rejection of ID samples from FFHQ, Celeb) and Power (correct rejection of moderate OOD (Met) and strong OOD (bedrooms, CBSD68, AFHQ) examples).
Figure 6: Samples from $x|y^{-}$ and $x|y^{+}$ for a correctly specified model (FFHQ) and a misspecified model (AFHQ), where $y$ is obtained by degrading an FFHQ image with increasing blur ($\sigma_{\kappa}=0.5,\;2,\;5$).
Figure 7: OOD detection on natural images, respectively for (a) $\sigma_{\kappa}=0.5$, (b) $\sigma_{\kappa}=2$ and (c) $\sigma_{\kappa}=5$. The dotted lines indicates the testing threshold ($95$-th percentile of the test statistic over the reference FFHQ subset).
Figure 8: Measurements $y$ and samples from $x|y^{-}$ and $x|y^{+}$ for the FFHQ-trained model, where $y$ is obtained by blurring a Met-Faces image.
4.3.2 MRI reconstruction
We now consider a single-coil MRI image reconstruction problem (see SM, Section 3). We use two diffusion priors from [44], which are trained on brain and knee images from the FastMRI dataset respectively [53, 26]. We consider the brain dataset as ID. We proceed similarly to the previous experiment and extract brain and knee scans from FastMRI to compose the ID and OOD datasets. For this experiment, we slightly increase $\alpha$ to $0.25$ to reduce the noise injected to $\mathbf{y}^{-}$, which allows reducing the number of noise realizations to $4$ and the number of steps to $10$. We define a reference dataset of $50$ brain scans to compute the $95$-th percentile of $\widehat{\Phi}^{2}_{y}$, and compose a test set of $50$ ID and $50$ OOD images. We set the measurement noise to $0.1$ in all experiments and consider an acceleration factor $R$ of $\times 4$ or $\times 8$ for the forward operator; increasing $R$ makes the estimation problem more challenging.
Figure 9: OOD detection on MRI scans at the acceleration factors (a) $R=4$ and (b) $R=8$. The dotted lines indicates the testing threshold, which corresponds to the $95$-th percentile of $\widehat{\Phi}^{2}_{y}$ (respectively $\widehat{\Phi}^{1}_{y}$) on the reference brain scan subset.
The values of the estimators $\widehat{\Phi}^{1}_{y}$ and $\widehat{\Phi}^{2}_{y}$ and for the brain-trained model are represented in Fig. 9 for $R=4$ and the more challenging case $R=8$. In both cases, we observe an excellent discrimination between ID and OOD data points for both estimators. This result can be explained by the fact that the brain images comprise few learnable features that can be transposed to knee images. The brain model mainly learns the complex structures (gyri) present on the surface of the brain, which are completely absent from knee scans, and tends to hallucinate these structures in knee reconstructions.
Figure 10: Samples from $x|y^{-}$ and $x|y^{+}$ for the brain-trained diffusion model on ID and OOD examples.
Moreover, to evaluate OOD detection accuracy, Tab. 3 reports the Type I error probability and testing power obtained with each estimator; we observe that they both achieve excellent performance. For completeness, we also report the results for single-shot model selection against a model trained on knee scans in SM, Section 3. Lastly, Fig. 10 shows examples of samples from $p(x|y^{-})$ and $p(x|y^{+})$ for ID and OOD cases. Once again, we observe that the OOD case exhibits substantial variability in perceptual details, largely hallucinated by the prior.
$R=4$ $R=8$
$\widehat{\Phi}^{2}_{y}$ $\widehat{\Phi}^{1}_{y}$ $\widehat{\Phi}^{2}_{y}$ $\widehat{\Phi}^{1}_{y}$
Type I Error 4% 6% 0% 4%
Power 100% 94% 100 % 96%
Table 3: Type I error rate (incorrect rejection of brain examples), Power (correct rejection of knee examples).
5 Discussion and conclusions
We introduced a Bayesian cross-validation framework for unsupervised model selection and misspecification testing in imaging inverse problems, with a focus on the objective comparison of likelihood functions and data-driven priors encoded by large-scale machine learning models. Leveraging data fission, the proposed method operates using only a single measurement, which is partitioned into two noisier measurements according to a parameter $\alpha$ that governs the amount of information reserved for model evaluation, as well as the trade-off between evaluation accuracy and computational cost. As the marginal likelihood, a gold standard for Bayesian model selection, is recovered in the limit as $\alpha\to 0$ and a specific choice of scoring rule, our approach can be viewed as a relaxation that sacrifices some accuracy for significant gains in efficiency. We propose two main scoring rules for evaluating Bayesian imaging models: a likelihood-based rule, well-suited for assessing likelihood functions, and a perceptual posterior-based rule, which effectively evaluates priors. Furthermore, we demonstrate the effectiveness of the proposed approach through a series of numerical experiments on image photographic deblurring and MRI reconstruction, showcasing its ability to compare likelihoods and image priors, as well as accurately diagnose prior misspecification in both mild and strong out-of-distribution settings.
Acknowledgements
This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011015754 and 2025-AD011015754R1 made by GENCI.
Reproducibility
The code to reproduce the experiments is available at https://github.com/aleph-group/Priors_selection, and the models and ground truth images can be downloaded from https://zenodo.org/records/17484892.
Appendix A Analysis in the Gaussian case
A.1 Derivation of the analytical formulas
We compute here the formula for $p(y^{+}|y^{-})$ for $y=x+e$, where $x\sim\mathcal{N}(0,\sigma_{x}^{2}I_{m})$ and $e\sim\mathcal{N}(0,\sigma^{2}I_{m})$. Recall that $y^{+}=y+\sqrt{\frac{\alpha}{1-\alpha}}w$ and $y^{-}=y-\sqrt{\frac{1-\alpha}{\alpha}}w$, with $w\sim\mathcal{N}(0,\sigma^{2}I_{m})$.
We have:
$p(y^{+}|y^{-})=\mathbb{E}_{x^{\prime}|y^{-}}\!\!\left[p(y^{+}|x^{\prime})\right]=\int p(y^{+}|x^{\prime})\frac{p(y^{-}|x^{\prime})p(x^{\prime})}{p(y^{-})}dx^{\prime}.$ (13)
We can thus write:
$p(y^{+}|y^{-})=$ $\int\frac{\alpha^{m/2}e^{-\frac{\alpha}{2\sigma^{2}}\|x^{\prime}-y^{-}\|^{2}}}{(2\pi)^{m/2}\sigma^{m}}\frac{(1-\alpha)^{m/2}e^{-\frac{1-\alpha}{2\sigma^{2}}\|x^{\prime}-y^{+}\|^{2}}}{(2\pi)^{m/2}\sigma^{m}}\frac{e^{-\frac{1}{2\sigma_{x}^{2}}\|x^{\prime}\|^{2}}}{(2\pi)^{m/2}\sigma_{x}^{m}}\frac{(2\pi)^{m/2}(\alpha\sigma_{x}^{2}+\sigma^{2})^{m/2}}{\alpha^{m/2}}e^{\frac{\alpha}{2(\alpha\sigma_{x}^{2}+\sigma^{2})}\|y^{-}\|^{2}}dx^{\prime}$ (14)
$=$ $\int\frac{\left[(1-\alpha)(\alpha\sigma_{x}^{2}+\sigma^{2})\right]^{m/2}}{(2\pi)^{m}\sigma^{2m}\sigma_{x}^{m}}e^{-\frac{\alpha}{2\sigma^{2}}\|x^{\prime}-y^{-}\|^{2}-\frac{1-\alpha}{2\sigma^{2}}\|x^{\prime}-y^{+}\|^{2}-\frac{1}{2\sigma_{x}^{2}}\|x^{\prime}\|^{2}+\frac{\alpha}{2(\alpha\sigma_{x}^{2}+\sigma^{2})}\|y^{-}\|^{2}}dx^{\prime}.$ (15)
The first part of the exponent can be factorized as:
$-\frac{1-\alpha}{2\sigma^{2}}\|x^{\prime}-y^{+}\|^{2}-\frac{\alpha}{2\sigma^{2}}\|x^{\prime}-y^{-}\|^{2}-\frac{1}{2\sigma_{x}^{2}}\|x^{\prime}\|^{2}=$ $-\frac{\sigma^{2}+\sigma_{x}^{2}}{2\sigma^{2}\sigma^{2}_{x}}\|x^{\prime}\|^{2}+\frac{1}{\sigma^{2}}x^{\prime}\cdot y-\frac{1-\alpha}{2\sigma^{2}}\|y^{+}\|^{2}-\frac{\alpha}{2\sigma^{2}}\|y^{-}\|^{2}$ (16)
$=$ $-\frac{\sigma^{2}+\sigma_{x}^{2}}{2\sigma^{2}\sigma^{2}_{x}}\|x^{\prime}-\frac{\sigma_{x}^{2}}{\sigma^{2}+\sigma^{2}_{x}}y\|^{2}+\frac{\sigma^{2}_{x}}{2\sigma^{2}(\sigma^{2}+\sigma_{x}^{2})}\|y\|^{2}$ (17)
$\quad-\frac{1-\alpha}{2\sigma^{2}}\|y^{+}\|^{2}-\frac{\alpha}{2\sigma^{2}}\|y^{-}\|^{2}.$
Integrating over $x^{\prime}$ yields:
$p(y^{+}|y^{-})=$ $\frac{((1-\alpha)(\alpha\sigma_{x}^{2}+\sigma^{2}))^{m/2}}{(2\pi)^{m}\sigma^{m}(\sigma^{2}+\sigma_{x}^{2})^{m/2}}e^{\frac{\sigma^{2}_{x}}{2\sigma^{2}(\sigma^{2}+\sigma_{x}^{2})}\|y\|^{2}-\frac{\alpha^{2}\sigma^{2}_{x}}{2\sigma^{2}(\alpha\sigma_{x}^{2}+\sigma^{2})}\|y^{-}\|^{2}-\frac{1-\alpha}{2\sigma^{2}}\|y^{+}\|^{2}}.$ (18)
Finally, expanding the norms in the exponential and refactoring leads to:
$p(y^{+}|y^{-})=\frac{((1-\alpha)(\alpha\sigma_{x}^{2}+\sigma^{2}))^{m/2}}{(2\pi)^{m}\sigma^{m}(\sigma^{2}+\sigma_{x}^{2})^{m/2}}e^{-\frac{1}{2(\alpha\sigma_{x}^{2}+\sigma^{2})}\left\|\sqrt{\frac{1-\alpha}{\sigma^{2}+\sigma_{x}^{2}}}\sigma y+\frac{\sqrt{\alpha(\sigma^{2}+\sigma_{x}^{2})}}{\sigma}w\right\|^{2}}.$ (19)
Note that as $\alpha\to 0$, we recover the density of $\mathbf{y}$, while the value vanishes to zero as $\alpha\to 1$.
Let $\widehat{p}(y^{+}|y^{-},\sigma_{x}^{2})$ be the approximation of $p(y^{+}|y^{-},\sigma_{x}^{2})$ computed by drawing from the posterior law $x|y^{-}$, following Eq. (12) of the main paper, either by using the analytical posterior law, or by simulating this distribution with the SK-ROCK algorithm [1]. Fig. 11 represents the relative error between the analytical value and the estimator as a function of the iterations for different values of $\alpha$ and dimensions of the target vector. Full lines correspond to Monte Carlo approximations of the posterior $x|y^{-}$, while the dotted lines are obtained by drawing from the analytical posterior. Both plots are obtained by averaging the error over $25$ samples of $\mathbf{w}$ for $\sigma_{x}=1$ and $\sigma=0.05$. The additional error caused by sampling with SK-ROCK seems negligible relative to the Monte Carlo integration error. Note that the convergence speed is slow, and the approximation error increases the relative error for a fixed number of iterations. Expectedly, the error also increases with the dimension $m$, as depicted in Fig. 12, where we plot the relative error as a function of $m$.
Figure 11: Relative log error between $\widehat{p}(y^{+},y^{-})$ and $p(y^{+},y^{-})$ as a function of the number of Monte Carlo integration steps NN, for different values of $\alpha$ and dimensions $m$. The full line is obtained by using the analytical posterior law, while the dotted line corresponds to SK-ROCK sampling.
Figure 12: Relative log error between $\widehat{p}(y^{+},y^{-})$ and $p(y^{+},y^{-})$ as a function of the dimension, using $N=50000$ MC steps and averaged over $25$ noise realizations, for $\alpha=0.5$.
Appendix B Kernel selection
B.1 Implementation details
$\kappa_{\mathcal{G}}(\sigma)$ $(x,y)\mapsto e^{-(x^{2}+y^{2})/2\sigma^{2}}$
$\kappa_{\mathcal{M}}(\sigma,\mu)$ $(x,y)\mapsto(\sigma^{2}(x^{2}+y^{2})/\mu+1)^{-(\mu/2+1)}$
$\kappa_{\mathcal{L}}(\sigma)$ $(x,y)\mapsto e^{\sigma(-|x|+|y|)}$
$\kappa_{\mathcal{U}}(s)$ $(x,y)\mapsto\mathds{1}_{x,y\leq s}$
Table 4: Unnormalized blurring kernels.
We give here some details on the implementation of the kernel selection experiment from the main paper. In all cases, we use the SK-ROCK algorithm to sample the posterior law, using $s=15$ inner iterations and the potential of the gradient step denoiser [19] as prior. We set the regularization parameter $\lambda$ to $110$ for every experiment when computing $\widehat{\Phi}_{y}^{1}$, based on a prior study of the reconstruction’s quality on a single observation. The standard deviation parameter of the denoiser is also fixed to $0.1$.
We re-use the same Markov chain to simulate both the prior and posterior laws in order to apply SAPG [50], adding respectively $15$ and $25$ thinning iterations before each posterior and prior sample. We initialize the algorithm with the reference value $110$, and perform $150$ SAPG iterations, generating approximately $6000$ samples in the process. While this number could be reduced by increasing the step size, this illustrates a limitation of SAPG, which requires careful per-application tuning to work best, especially when a good first estimate is unavailable. In contrast, we use a single chain to generate the samples in our method, using $20$ thinning iterations before swapping to a new noise realization, for a total of approximately $1200$ sampling steps. Note also that, as we did not tune the regularization parameter for our method, the prior’s misspecification is higher. We use a single V100 16GB GPU to process a single image.
Note that we used a FFT-based blur operator, which involves the application of circular padding. In order to avoid a potential bias due to this padding, we ignore the padding pixels when computing each metric (i.e. use ”valid” padding). The amount of pixels removed is based on the span of the largest convolutional kernel. We used an implementation based on the Deepinv library [48] for the forward operator. The analytical definition of each kernel is given in Tab. 4.
B.2 Numerical results
  GT Test $\kappa_{\mathcal{G}}(2)$ $\kappa_{\mathcal{M}}(0.5,1)$ $\kappa_{\mathcal{L}}(0.4)$ $\kappa_{\mathcal{U}}(3)$ $\kappa_{\mathcal{G}}(2.5)$
I1 I2 I3 I1 I2 I3 I1 I2 I3 I1 I2 I3 I1 I2 I3
$\kappa_{\mathcal{G}}(2)$ 45.72 58.83 56.87 46.01 60.50 57.31 51.79 67.34 59.42 49.90 65.49 58.44 48.07 61.60 59.41
$\kappa_{\mathcal{M}}(0.5,1)$ 46.50 62.43 53.96 44.32 56.76 53.64 45.55 58.81 54.48 50.10 67.69 54.80 48.14 60.16 54.95
$\kappa_{\mathcal{L}}(0.4)$ 41.09 55.26 45.89 39.63 54.04 44.96 38.55 52.97 44.73 42.10 56.78 45.88 39.86 54.73 45.10
$\kappa_{\mathcal{U}}(3)$ 46.35 60.02 53.26 48.37 62.59 53.75 51.55 68.36 55.09 45.10 57.07 53.46 47.03 58.07 54.36
$\kappa_{\mathcal{G}}(2.5)$ 40.31 53.05 46.42 40.16 53.03 46.14 41.02 54.02 46.13 40.56 53.41 46.44 39.39 52.46 46.29
Table 5: Value of $\widehat{\Phi}^{1}_{y}-1100$ for the three test images for different ground truth blurring kernel (rows), computed using 10 noise realizations with 100 steps each and $\alpha=0.5$. The best values for each row are highlighted in bold font, with a mean accuracy of $86.7\%$ over the $15$ experiments.
 GT Test $\kappa_{\mathcal{G}}(2)$ $\kappa_{\mathcal{M}}(0.5,1)$ $\kappa_{\mathcal{L}}(0.4)$ $\kappa_{\mathcal{U}}(3)$ $\kappa_{\mathcal{G}}(2.5)$
$\kappa_{\mathcal{G}}(2)$ 13.808 14.603 19.515 17.942 16.361
$\kappa_{\mathcal{M}}(0.5,1)$ 14.295 11.574 12.947 17.534 14.416
$\kappa_{\mathcal{L}}(0.4)$ 7.416 6.210 5.414 8.252 6.566
$\kappa_{\mathcal{U}}(3)$ 13.213 14.904 18.335 11.875 13.155
$\kappa_{\mathcal{G}}(2.5)$ 6.592 6.443 7.055 6.802 6.045
Table 6: Values of $\widehat{\Phi}^{1}_{y}(\mathcal{M})-1100$ averaged over three test images for different ground truth blurring kernel (rows), computed using 10 noise realizations with 100 steps each and $\alpha=0.5$.
 GT Test $\kappa_{\mathcal{G}}(2)$ $\kappa_{\mathcal{M}}(0.5,1)$ $\kappa_{\mathcal{L}}(0.4)$ $\kappa_{\mathcal{U}}(3)$ $\kappa_{\mathcal{G}}(2.5)$
$\kappa_{\mathcal{G}}(2)$ 14.936 14.955 19.647 22.448 22.378
$\kappa_{\mathcal{M}}(0.5,1)$ 17.353 12.396 17.230 22.374 20.946
$\kappa_{\mathcal{L}}(0.4)$ 12.418 11.050 10.991 15.277 15.731
$\kappa_{\mathcal{U}}(3)$ 14.515 16.809 16.773 17.737 18.355
$\kappa_{\mathcal{G}}(2.5)$ 10.825 9.085 10.620 12.235 12.927
Table 7: Values of $\|\kappa\ast\widehat{x}_{\kappa}-y_{\kappa_{\operatorname{GT}}}\|_{2}^{2}$ - 560 averaged over three test images, where $\widehat{x}_{\kappa}$ denotes the approximate MAP reconstruction using the tested blurring kernel $\kappa$ for the forward model.
We report in this section the numerical results from the kernel selection experiments. Tab. 5 displays the values of $\widehat{\Phi}_{y}^{1}(\mathcal{M})$. Each row corresponds to different measurements generated using different blurring kernels, while the columns correspond to the tested kernels. I1, I2, I3 denote the three test images depicted in the main paper. Tab. 6 gives the average value of the estimator over I1, I2, I3. The estimator fails to select the correct kernel in only two out of fifteen cases when using a single observation and selects the correct kernel in every case when averaging over the three test images. The values of the estimator are quite close with the number of samples used. The single-shot performance might still be improved by increasing the samples in the Monte Carlo approximation. Tab. 7 gives the average (unnormalized) of the MAP reconstruction, obtained after tuning the regularization parameter by applying SAPG. Even in the few-shot setting, we only reach $60\%$ accuracy.
Figure 13: Distributions of the values taken by $\widehat{\Phi}^{1}_{y}(\mathcal{M})$ and $\widehat{\Phi}^{2}_{y}(\mathcal{M})$ over the FFHQ subset, for the FFHQ and AFHQ-trained models, at $\sigma_{\kappa}=0.5,2,5$ and $\alpha=0.1$.
Figure 14: Samples from $x|y^{-}$ and $x|y^{+}$ for the FFHQ and AFHQ-trained models, where $y$ is obtained by blurring a FFHQ image with $\sigma_{\kappa}=5$.
Appendix C Misspecification detection
C.1 Deblurring of natural images
We provide here some figures to further illustrate the observations made in Section 4.3.1 of the main paper. Fig. 13 represents the distributions of $\widehat{\Phi}^{1}_{y}(\mathcal{M})$ and $\widehat{\Phi}^{2}_{y}(\mathcal{M})$ for the FFHQ and AFHQ-trained models at different blur levels over the FFHQ subset. As the blur level increases, the values of $\widehat{\Phi}^{2}_{y}(\mathcal{M})$ spread out, while the distribution of $\widehat{\Phi}^{1}_{y}(\mathcal{M})$ becomes sharper. Indeed, at higher blur values, the inter-sample differences are small in the measurement space, while the sample variety increases.
Fig. 15 depicts the distributions of $\widehat{\Phi}^{1}_{y}(\mathcal{M})$ and $\widehat{\Phi}^{2}_{y}(\mathcal{M})$ for the FFHQ and AFHQ-trained models for different values of $\alpha$, at $\sigma_{\kappa}=0.5$ and $\sigma_{\kappa}=5$. The perceptual variance of the samples greatly increases when we let $\alpha$ reduce for the OOD model, while it is less affected for the ID model. Fig. 16 shows that a poor choice of $\alpha$ translates into increased statistical error rates when using $\widehat{\Phi}^{2}_{y}(\mathcal{M})$ for model selection. Indeed, when $\alpha$ is close to $0.5$, the noise quantity imbalance between $y^{+}$ and $y^{-}$ vanishes, and the perceptual variance between samples is reduced, rendering the task of detecting OOD images from sample variance less effective. Note that, when the amount of information available in the measurements is low, the perceptual variance of the samples is high even for ID images. The variations in specific details of the samples, such as a mouth being open or closed, can cause the test to fail in extreme cases. Fig. 14 shows such an example, where $\widehat{\Phi}^{2}_{y}(\mathcal{M})$ is slightly higher for the FFHQ-trained model. Note, however, that $\widehat{\Phi}^{1}_{y}(\mathcal{M})$ chooses the correct model in this case.
Finally, a concern can be raised by observing the low convergence speed of the estimator in the Gaussian analytical case (see Fig. 11). The experiments on natural images show that fine convergence is not required in order to have accurate model selection or misspecification detection. Fig. 17 shows the value of the estimator $\widehat{\Phi}^{2}_{y}(\mathcal{M})$ as a function of the number of $x|y^{-}$ samples. We can observe that, even though the estimator has not converged, the variation w.r.t. new iterations is negligible. However, in border cases where the information available is low, such as the case depicted in Fig. 14, adding some iterations might improve results.
(a) $\sigma_{\kappa}=0.5$
(b) $\sigma_{\kappa}=5$
Figure 15: Distributions of the values taken by $\widehat{\Phi}^{1}_{y}(\mathcal{M})$ and $\widehat{\Phi}^{2}_{y}(\mathcal{M})$ over the FFHQ subset, for the FFHQ and AFHQ-trained models, at $\alpha=0.1,0.3,0.5$.
Figure 16: Type 1 error rate on ID images, i.e. rejection rate on FFHQ, Celeb test sets, and type 2 error rates, i.e. acceptance rate for moderately OOD (Met-Faces) and strongly OOD (Bedrooms, CBSD68, AFHQ) images, as a function of $\alpha$ for the deblurring problem with $\sigma_{\kappa}=0.5$.
Figure 17: $\widehat{\Phi}^{2}_{y}(\mathcal{M})$ as a function of the number of steps NN, at a fixed number of noise realizations $K=10$, for a single Celeb-Faces image (a) and for each image of the test dataset (b).
$R=4$ $R=8$
$\widehat{\Phi}^{2}_{y}$ $\widehat{\Phi}^{1}_{y}$ $\widehat{\Phi}^{2}_{y}$ $\widehat{\Phi}^{1}_{y}$
Brain 86% 60% 92% 76%
Knee 88% 96% 82% 96%
Table 8: Accuracy of model selection on the brain and knee scan datasets using $\widehat{\Phi}^{2}_{y}$ and $\widehat{\Phi}^{1}_{y}$.
C.2 MRI reconstruction
We give here additional illustrations and details for Section 4.3.2 of the main paper. The forward model for the single-coil accelerated MRI problem writes:
$y=M\mathcal{F}x,$ (20)
where $\mathcal{F}$ denotes the 2D Fourier transform, and $M$ is the sub-sampling operator that applies a mask to the Fourier observations. For simplicity, we do not consider coil sensitivity matrices. In practical experiments, the observations have a fixed under-sampling at low frequencies, and random Gaussian under-sampling at high frequencies. As in the previous section, we use an implementation based on Deepinv [48] to apply the DiffPIR algorithm.
We perform single-shot model selection by computing the estimators on both datasets using both models. The accuracy of each estimator’s prediction is reported in Tab. 8. $\widehat{\Phi}^{2}_{y}(\mathcal{M})$ performs better than $\widehat{\Phi}^{1}_{y}(\mathcal{M})$ when comparing the models on brain images, but fares worse on knee images. This can be partially explained by the fact that the knee model seems slightly under-trained and sometimes produces low-quality knee samples. Fig. 18a displays an image for which $\widehat{\Phi}^{2}_{y}(\mathcal{M})$ incorrectly favors the brain model over the knee model, while $\widehat{\Phi}^{1}_{y}(\mathcal{M})$ selects the correct model. The brain-trained model hallucinates brain features in its samples, but the perceptual quality of these reconstructions still ranks higher than the knee-trained model’s samples. Fig. 18b gives an example of a brain scan for which both estimators select the correct model. Some of the brain’s features are recovered by the knee-trained model in samples from $x|y^{+}$, but are lost in samples from $x|y^{-}$ due to the added noise. The overall lower performance of $\widehat{\Phi}^{2}_{y}(\mathcal{M})$ can also be explained by the fact that the perceptual metric was trained on natural images, and fine-tuning this metric on MRI images might improve results.
(a) Knee scan
(b) Brain scan
Figure 18: Samples from $x|y^{-}$ and $x|y^{+}$ for the brain and knee-trained models, where $y$ is an under-sampled scan with $R=4$.